OpenStudy (anonymous):

simplify the expression (7i)(4i)

OpenStudy (anonymous):

-28 i*i = -1 7*4 = 28

OpenStudy (anonymous):

(7*4)(i*i)=28*(-1)=-28

OpenStudy (anonymous):

thanks both answers are the best answers